### v1

@@ 3.1 Well-posedness of the Stokes-Darcy problem
Here the use of standard $L^{2}$ simplifies the implementation. However, the space introduces a new challenge: the inf-sup constant depends on the physical parameters, which is discussed in the following theorem.
@@
For the NE and EN cases, the system is well-posed on the space $\widetilde{\mathbf{W}}$. Then the Stokes-Darcy operator $\mathcal{A}$ in (2.7) is an isomorphism mapping $\widetilde{\mathbf{W}}$ to $\widetilde{\mathbf{W}}^{\prime}$ such that $\|\mathcal{A}\|_{\mathcal{L}\left(\widetilde{\mathbf{W}},\widetilde{\mathbf{W}}^{\prime}\right)}\leq C_{1}$ and $\left\|\mathcal{A}^{-1}\right\|_{\mathcal{L}\left(\widetilde{\mathbf{W}}^{\prime},\widetilde{\mathbf{W}}\right)}\leq$ $C_{2}^{-1}$, where $C_{1}$ and $C_{2}$ depend on the physical parameters $\mu$ and ${K}$.
@@
(3.20a) $2\mu(\bm{\epsilon}(\mathbf{u}_{S}),\bm{\epsilon}(\mathbf{v}_{S}))_{S}+\beta_{\tau}(T_{t}\mathbf{u}_{S},T_{t}\mathbf{v}_{S})_{\Gamma}+(\nabla\cdot\mathbf{v}_{S},p_{S})_{S}+(T_{S}\mathbf{v}_{S},\lambda_{S})_{\Gamma}=0,$ $\quad\forall\,\mathbf{v}_{S}\in\mathbf{V}_{S},$
@@
(3.23a) $K^{-1}(\mathbf{u}_{D},\mathbf{v}_{D})_{D}+(\nabla\cdot\mathbf{v}_{D},p_{D})_{D}+(T_{D}\mathbf{v}_{D},\lambda_{D})_{\Gamma}=0,$ $\quad\forall\,\mathbf{v}_{D}\in\mathbf{V}_{D},$
@@ 3.2 Preconditioning
Building on the well-posedness results from Theorem 3.7 and Theorem 3.8, and leveraging the operator preconditioning framework [30], the canonical preconditioner for the coupled Stokes-Darcy problem is given by
@@
[…]
@@
By the Courant–Fischer–Weyl min-max principle, Theorem 3.7 and Theorem 3.11 indicate that we can divide $\sigma(\mathcal{BA})$, the spectrum of $\mathcal{BA}$, into two parts: $\{\lambda_{0}\}$ and $\sigma_{1}(\mathcal{BA})$. The eigenvalue $\lambda_{0}$, which has the smallest absolute value, depends on the physical parameters $\mu$ and ${K}$. On the other hand, $\sigma_{1}(\mathcal{BA})\subset[-C,-\gamma]\cup[\gamma,C]$ contains the remaining eigenvalues that are bounded with respect to the parameters. Based on the theory of the MINRES method [42], the convergence rate can be split into two parts; 1) a period of stagnation related to the removal of the eigencomponents associated with the smallest eigenvalue and 2) the convergence rate predicted by $(\kappa_{\mathrm{eff}}(\mathcal{BA})-1)/(\kappa_{\mathrm{eff}}(\mathcal{BA})+1)$, which is parameter-robust.
@@ 3.3 Robustness of discrete preconditioner
(3.26) $\mathbf{V}_{S,h}$ =$=\left\{\mathbf{v}\in\mathbf{H}^{1}(\Omega_{S})\,|\,\mathbf{v}|_{K}\in[\mathbb{P}_{2}(K)]^{d}\ \ \forall\,K\in\mathcal{T}_{S,h}\right\},$
$Q_{S,h}$ =$=\left\{q\in{H}^{1}(\Omega_{S})\,|\,{q}|_{K}\in\mathbb{P}_{1}(K)\ \ \forall\,K\in\mathcal{T}_{S,h}\right\},$
@@
(3.27) $\mathbf{V}_{D,h}$ =$=\left\{\mathbf{v}\in\mathbf{H}(\text{div},\Omega_{D})\,|\,\mathbf{v}|_{K}\in\mathbb{RT}_{0}(K)\ \ \forall\,K\in\mathcal{T}_{D,h}\right\},$
$Q_{D,h}$ =$=\left\{q\in{L}^{2}(\Omega_{D})\,|\,{q}|_{K}\in\mathbb{P}_{0}(K)\ \ \forall\,K\in\mathcal{T}_{D,h}\right\},$
@@
$\Lambda_{h}$ =$=\left\{\lambda\in L^{2}(\Gamma)\,|\,\lambda|_{F}\in\mathbb{P}_{0}(F)\ \ \forall\,F\in\Gamma_{h}\right\}.$
@@
We implemented the above discretization using the FEniCS finite element library [28] and its extension FEniCSii [26] to handle the coupling at the interface. Since the spaces $\mathbf{V}_{S,h}$ and $\mathbf{V}_{D,h}$ are $H^{1}$- and $H(\text{div})$-conforming, respectively, the discrete operators $\mathcal{A}_{h}$ and $\mathcal{B}_{h}$ are obtained by applying $\mathcal{A}$ and $\mathcal{B}$ to $\mathbf{W}_{h}\subset\mathbf{W}$, respectively. We note that the discrete representation of the multiplier preconditioner $S$ in (3.25) is computed using a spectral/Fourier representation, see, for example, [5].
@@
[…]
@@
To assess robustness of the preconditioner, we typically consider the spectral condition number of the discrete preconditioned problem. The associated generalized eigenvalue problem $\mathcal{A}_{h}x=\lambda\mathcal{B}^{-1}_{h}x$ is solved iteratively using SLEPc [24] for the largest eigenvalue in magnitude and the smallest (for $\kappa$) or the second smallest (for $\kappa_{\text{eff}}$) eigenvalues in magnitude. Additionally, we will discuss performance of the preconditioners in terms of boundedness of the MINRES iterations. In this context, we apply the exact preconditioner with the diagonal blocks inverted using LU decomposition. However, we remark that all the blocks can be approximated by scalable optimal methods. Specifically, the fractional interface preconditioners can be efficiently implemented, for example, through rational approximations [7, 1].
@@ 4.1 Convergence analysis
where the $n\times n$ matrix $\mathcal{A}$ is symmetric and indefinite. Given the initial guess $x_{0}$ to the solution of (4.1), and the corresponding initial residual $r_{0}=b-\mathcal{A}x_{0}$, the Krylov subspace of dimension $m$ defined by $\mathcal{A}$ and $r_{0}$ is given by $\mathcal{K}_{m}(\mathcal{A},r_{0})=\operatorname{span}\{r_{0},\mathcal{A}r_{0},...,\mathcal{A}^{m-1}r_{0}\}$. At the $m$-th step, an approximation to the solution of (4.1), $x_{m}$, can be obtained in $x_{0}+\mathcal{K}_{m}(\mathcal{A},r_{0})$, by imposing some additional conditions. The residual $r_{m}$ is related to the initial residual as $r_{m}=p_{m}(\mathcal{A})r_{0},$ where the residual polynomial $p_{m}\in\mathbb{P}_{m}$, the set of polynomials of degree at most $m$ such that $p_{m}(0)=1$. Let $\Lambda(\mathcal{A})=\{\lambda_{1},\lambda_{2},...,\lambda_{n}\}$ be the set of eigenvalues and $\{v_{1},v_{2},...,v_{n}\}$ be the corresponding orthonormal eigenvectors of $\mathcal{A}$, with the eigenvalues ordered increasingly in magnitude. Then we have the following standard bound [42]
@@
A more specific bound was developed in [21, Section 3.1] for the case where $\Lambda(\mathcal{A})\subset[a,b]\cup[c,d]$, $a<b<0<c<d$. Under the constraint that $|b-a|=|d-c|$, using an appropriate transformation of the intervals and bounds on Chebychev polynomials, the following estimate holds
@@
Following [42], let us now assume that $\lambda_{1}$ is close to zero, and the remaining eigenvalues belong to $[a,b]\cup[c,d]$. Let $r_{0}=\sum\limits_{k=1}^{n}\alpha_{k}v_{k}$, where $\alpha_{k}=(r_{0},v_{k}),k=1,...,n$. Then, we have
@@
Let $\bar{q}_{j}$ be the MINRES polynomial in $\mathcal{K}_{j}(\mathcal{A},\bar{r}_{0})$, then $\bar{r}_{j}=\bar{q}_{j}(\mathcal{A})\bar{r}_{0}=\sum_{k=2}^{n}\bar{q}_{j}(\lambda_{k})p_{m}(\lambda_{k})\alpha_{k}v_{k}$. Let $\theta_{1}^{(m)}$ be the harmonic Ritz value (see [42, 32, 35]) closet to $\lambda_{1}$ in $\mathcal{K}_{m}(\mathcal{A},r_{0})$, and define the polynomial
@@
In the next section, we employ the deflation technique (e.g. [12]) to remove the plateau region and accelerate the MINRES convergence.
@@ 4.2 Deflation technique
where $P_{W}$ is defined by an approximations of the slow-to-converge eigenmodes and $\gamma>0$ is a problem-dependent scaling parameter.
@@
We remark that the scaling parameter $\gamma$ reflects the dependence of the inf-sup constant in Theorem 3.8. Since the near-kernel is one dimensional, the matrix representation of $P_{W}$ is the vector representation of $w$ in $W_{h}$. As as result, computing the small $(1\times 1)$ matrix $(P_{W}^{T}\gamma\mathcal{B}^{-1}P_{W})$ and its inverse can be done efficiently with optimal complexity.
@@
We study the Stokes-Darcy problem in geometries with multiple disconnected porous subdomains, This setup is inspired by [8] in the context of tissue bioengineering, particularly in the design of scaffolds modeled as porous structures. As a representative example, we consider a channel domain containing $12$ Darcy subdomains, denoted as $\Omega^{i}_{D}$ for $i=1,\dots,12$. Each subdomain has a hexagonal shape with a radius of $1$ and is arranged such that the free-flow region forms channels of width $0.2$, as illustrated in Figure 11. We impose a no-slip condition, $\mathbf{u}_{S}=0$, on the top and bottom walls of the channel, while the flow is driven by a traction boundary condition of the form $-\mathbf{\sigma}(\mathbf{u}_{S},p_{S})\cdot\mathbf{n}_{S}=p\mathbf{n}S$, where $p=1$ on the left edge and $p=0$ on the right edge. Although the coupled problem remains well-posed, the absence of a Neumann boundary on the Darcy subdomains ($\lvert\Gamma^{N}_{D}\rvert=0$) poses a preconditioning challenge similar to the NE case in Theorem 3.8. To illustrate this issue, we fix $\mu=3$ and $\alpha_{\text{BJS}}=0.5$ while varying $K\in\left\{1,100\right\}$. The resulting flow field for $K=100$ is depicted in Figure 11.

### v2

@@ 3.1 Well-posedness of the Stokes-Darcy problem
Here the use of standard $L^{2}$ simplifies implementation. However, the space introduces a new challenge: inf-sup constant depends on physical parameters, which is discussed in the following theorem.
@@
For the NE and EN cases, the system is well-posed on the space $\widetilde{\mathbf{W}}$. Then the Stokes-Darcy operator $\mathcal{A}$ in (2.7) is an isomorphism mapping $\widetilde{\mathbf{W}}$ to $\widetilde{\mathbf{W}}^{\prime}$ such that $\|\mathcal{A}\|_{\mathcal{L}\left(\widetilde{\mathbf{W}},\widetilde{\mathbf{W}}^{\prime}\right)}\leq C_{1}$ and $\left\|\mathcal{A}^{-1}\right\|_{\mathcal{L}\left(\widetilde{\mathbf{W}}^{\prime},\widetilde{\mathbf{W}}\right)}\leq$ $C_{2}^{-1}$, where $C_{1}$ and $C_{2}$ depend on the physical parameters $\mu$ and K{K}.
@@
(3.21a) $2\mu(\bm{\epsilon}(\mathbf{u}_{S}),\bm{\epsilon}(\mathbf{v}_{S}))_{S}+\beta_{\tau}(T_{t}\mathbf{u}_{S},T_{t}\mathbf{v}_{S})_{\Gamma}-(\nabla\cdot\mathbf{v}_{S},p_{S})_{S}+(T_{S}\mathbf{v}_{S},\lambda_{S})_{\Gamma}=0,$ $\quad\forall\,\mathbf{v}_{S}\in\mathbf{V}_{S},$
@@
(3.24a) $K^{-1}(\mathbf{u}_{D},\mathbf{v}_{D})_{D}-(\nabla\cdot\mathbf{v}_{D},p_{D})_{D}+(T_{D}\mathbf{v}_{D},\lambda_{D})_{\Gamma}=0,$ $\quad\forall\,\mathbf{v}_{D}\in\mathbf{V}_{D},$
@@ 3.2 Preconditioning
Building on the well-posedness results from Theorem 3.7 and Theorem 3.8, and leveraging the operator preconditioning framework [Kent2011Preconditioning], the canonical preconditioner for the coupled Stokes-Darcy problem is given by
@@
(3.26) $(\text{NN})\,\,$ $S=\mu^{-1}(-\Delta+I)^{-1/2}+{K}(-\Delta+I)_{00}^{1/2},$
$(\text{EE})\,\,$ $S=\mu^{-1}(-\Delta+I)_{00}^{-1/2}+{K}(-\Delta+I)^{1/2},$
@@
By the Courant–Fischer–Weyl min-max principle, Theorem 3.7 and Theorem 3.11 indicate that we can divide $\sigma(\mathcal{BA})$, the spectrum of $\mathcal{BA}$, into two parts: $\{\lambda_{0}\}$ and $\sigma_{1}(\mathcal{BA})$. The eigenvalue $\lambda_{0}$, which has the smallest absolute value, depends on the physical parameters $\mu$ and K{K}. On the other hand, $\sigma_{1}(\mathcal{BA})\subset[-C,-\gamma]\cup[\gamma,C]$ contains the remaining eigenvalues that are bounded with respect to the parameters. Based on the theory of the MINRES method [simoncini2013superlinear], the convergence rate can be split into two parts; 1) a period of stagnation related to the removal of the eigencomponents associated with the smallest eigenvalue and 2) the convergence rate predicted by $(\kappa_{\mathrm{eff}}(\mathcal{BA})-1)/(\kappa_{\mathrm{eff}}(\mathcal{BA})+1)$, which is parameter-robust.
@@ 3.3 Robustness of discrete preconditioner
(3.27) $\mathbf{V}_{S,h}$ $=\left\{\mathbf{v}\in\mathbf{H}^{1}(\Omega_{S})\,|\,\mathbf{v}|_{K}\in[\mathbb{P}_{2}(K)]^{d}\ \ \forall\,K\in\mathcal{T}_{S,h}\right\},$
$Q_{S,h}$ $=\left\{q\in{H}^{1}(\Omega_{S})\,|\,{q}|_{K}\in\mathbb{P}_{1}(K)\ \ \forall\,K\in\mathcal{T}_{S,h}\right\},$
@@
(3.28) $\mathbf{V}_{D,h}$ $=\left\{\mathbf{v}\in\mathbf{H}(\text{div},\Omega_{D})\,|\,\mathbf{v}|_{K}\in\mathbb{RT}_{0}(K)\ \ \forall\,K\in\mathcal{T}_{D,h}\right\},$
$Q_{D,h}$ $=\left\{q\in{L}^{2}(\Omega_{D})\,|\,{q}|_{K}\in\mathbb{P}_{0}(K)\ \ \forall\,K\in\mathcal{T}_{D,h}\right\}.$
@@
$\Lambda_{h}$ $=\left\{\lambda\in L^{2}(\Gamma)\,|\,\lambda|_{F}\in\mathbb{P}_{0}(F)\ \ \forall\,F\in\Gamma_{h}\right\}.$
@@
We implemented the above discretization using the FEniCS finite element library [logg2012automated] and its extension FEniCSii [kuchta2020assembly] to handle the coupling at the interface. Since the spaces $\mathbf{V}_{S,h}$ and $\mathbf{V}_{D,h}$ are $H^{1}$- and $H(\text{div})$-conforming, respectively, the discrete operators $\mathcal{A}_{h}$ and $\mathcal{B}_{h}$ are obtained by applying $\mathcal{A}$ and $\mathcal{B}$ to $\mathbf{W}_{h}\subset\mathbf{W}$, respectively. We note that the discrete representation of the multiplier preconditioner $S$ in (3.26) is computed using a spectral/Fourier representation, see, for example, [boon2022robust].
@@
The Darcy flux $\mathbf{u}_{D}$ is computed from its definition based on the pressure $p_{D}$. We note that the solution (3.29)-(3.30) requires additional source terms on the right-hand side of the coupling conditions. This data, alone with the volume source terms $\bm{f}_{S}$ and $g_{D}$ is set according to the material parameters given as $\mu=3$, K=1{K}=1, $\alpha_{\text{BJS}}=0.5$. Finally, we prescribe mixed boundary conditions on the boundaries
@@
To assess robustness of the preconditioner, we typically consider the spectral condition number of the discrete preconditioned problem. The associated generalized eigenvalue problem $\mathcal{A}_{h}x=\lambda\mathcal{B}^{-1}_{h}x$ is solved iteratively using SLEPc [Hernandez:2005:SSF] for the largest eigenvalue in magnitude and the smallest (for $\kappa$) or the second smallest (for $\kappa_{\text{eff}}$) eigenvalues in magnitude. Additionally, we will discuss performance of the preconditioners in terms of boundedness of the MINRES iterations. In this context, we apply the exact preconditioner with the diagonal blocks inverted using LU decomposition. However, we remark that all the blocks can be approximated by scalable optimal methods. Specifically, the fractional interface preconditioners can be efficiently implemented, for example, through rational approximations [budivsa2022rational, adler2024improving].
@@ 4.1 Convergence analysis
where the $n\times n$ matrix $\mathcal{A}$ is symmetric and indefinite. Given the initial guess $x_{0}$ to the solution of (4.1), and the corresponding initial residual $r_{0}=b-\mathcal{A}x_{0}$, the Krylov subspace of dimension $m$ defined by $\mathcal{A}$ and $r_{0}$ is given by $\mathcal{K}_{m}(\mathcal{A},r_{0})=\operatorname{span}\{r_{0},\mathcal{A}r_{0},...,\mathcal{A}^{m-1}r_{0}\}$. At the $m$-th step, an approximation to the solution of (4.1), $x_{m}$, can be obtained in $x_{0}+\mathcal{K}_{m}(\mathcal{A},r_{0})$, by imposing some additional conditions. The residual $r_{m}$ is related to the initial residual as $r_{m}=p_{m}(\mathcal{A})r_{0},$ where the residual polynomial $p_{m}\in\mathbb{P}_{m}$, the set of polynomials of degree at most $m$ such that $p_{m}(0)=1$. Let $\Lambda(\mathcal{A})=\{\lambda_{1},\lambda_{2},...,\lambda_{n}\}$ be the set of eigenvalues and $\{v_{1},v_{2},...,v_{n}\}$ be the corresponding orthonormal eigenvectors of $\mathcal{A}$, with the eigenvalues ordered increasingly in magnitude. Then we have the following standard bound [simoncini2013superlinear]
@@
A more specific bound was developed in [Greenbaum1997Iterative, Section 3.1] for the case where $\Lambda(\mathcal{A})\subset[a,b]\cup[c,d]$, $a<b<0<c<d$. Under the constraint that |ba|=|dc||b-a|=|d-c|, using an appropriate transformation of the intervals and bounds on Chebychev polynomials, the following estimate holds
@@
Following [simoncini2013superlinear], let us now assume that $\lambda_{1}$ is close to zero, and the remaining eigenvalues belong to $a,b]\cup[c,d]$. Let $r_{0}=\sum\limits_{k=1}^{n}\alpha_{k}v_{k}$, where $\alpha_{k}=(r_{0},v_{k}),k=1,...,n$. Then, we have
@@
Let $\bar{q}_{j}$ be the MINRES polynomial in $\mathcal{K}_{j}(\mathcal{A},\bar{r}_{0})$, then $\bar{r}_{j}=\bar{q}_{j}(\mathcal{A})\bar{r}_{0}=\sum_{k=2}^{n}\bar{q}_{j}(\lambda_{k})p_{m}(\lambda_{k})\alpha_{k}v_{k}$. Let $\theta_{1}^{(m)}$ be the harmonic Ritz value (see [simoncini2013superlinear, morgan1991computing, paige1995approximate]) closet to $\lambda_{1}$ in $\mathcal{K}_{m}(\mathcal{A},r_{0})$, and define the polynomial
@@
In the next section, we employ the deflation technique (e.g. [coulaud2013deflation]) to remove the plateau region and accelerate the MINRES convergence.
@@ 4.2 Deflation technique
where $P_{W}$ is a projection operator onto the subspace spanned by approximations of the slow-to-converge eigenmodes and $\gamma>0$ is a problem-dependent scaling parameter.
@@
We remark that the scaling parameter $\gamma$ reflects the dependence of the inf-sup constant in Theorem 3.8. Since the near-kernel is one dimensional, the matrix representation of $P_{W}$ is a matrix consisting of a single column, where the column vector represents the function $w$ in $W_{h}$. As as result, computing the small $(1\times 1)$ matrix $(P_{W}^{T}\gamma\mathcal{B}^{-1}P_{W})$ and its inverse can be done efficiently with optimal complexity.
@@
We study the Stokes-Darcy problem in geometries with multiple disconnected porous subdomains, This setup is inspired by [bukavc2023analysis] in the context of tissue bioengineering, particularly in the design of scaffolds modeled as porous structures. As a representative example, we consider a channel domain containing $12$ Darcy subdomains, denoted as $\Omega^{i}_{D}$ for $i=1,\dots,12$. Each subdomain has a hexagonal shape with a radius of $1$ and is arranged such that the free-flow region forms channels of width $0.2$, as illustrated in Figure 11. We impose a no-slip condition, $\mathbf{u}_{S}=0$, on the top and bottom walls of the channel, while the flow is driven by a traction boundary condition of the form $\mathbf{\sigma}(\mathbf{u}_{S},p_{S})\cdot\mathbf{n}_{S}=p\mathbf{n}S$, where $p=1$ on the left edge and $p=0$ on the right edge. Although the coupled problem remains well-posed, the absence of a Neumann boundary on the Darcy subdomains ($\lvert\Gamma^{N}_{D}\rvert=0$) poses a preconditioning challenge similar to the NE case in Theorem 3.8. To illustrate this issue, we fix $\mu=3$ and $\alpha_{\text{BJS}}=0.5$ while varying $K\in\left\{1,100\right\}$. The resulting flow field for $K=100$ is depicted in Figure 11.